# Should Teleoperation Be like Driving in a Car? Comparison of Teleoperation HMIs

### v1

@@ II-A Taxonomy
A teleoperation system comprises three components, as shown in Fig. 1:
@@ II-C Separate Path and Velocity Specification
Initial research into control via waypoints sent to a real vehicle was conducted by Kay [31] in a study with 19 participants. Vehicle velocity was controlled by a safety driver in the vehicle The aim of the study was to investigate the functionality and performance of the waypoint specification under the aspect of latency and limited image data by measuring Key Performance Indicators like retries, total time or mean number of sent points. Effort in improving the RO workload is missing.
@@ III-A Design
To decide on an HMI design for separate path and velocity input, functionalities were collected in a morphological box and discussed with experts from the Institute of Automotive Technology at the Technical University of Munich. The discussion resulted in an HMI design that contains the components shown in Fig. 3.
To enable the RO to set a valid path, the vehicle’s surroundings are represented on two screens showing a camera image and a map where the waypoints are set. Thereby, the monitor and laptop only serve as display devices. First, the RO specifies individual waypoints on the map by clicking the mouse in HMI variant 2 or using the stylus on the touchscreen in 3. The waypoints are automatically combined in the clicked sequence to form a driving path. During the interaction, a new waypoint is already selected via a hover function, so the RO only has to set the point with one click. Furthermore, the user can edit the path by clicking, holding and moving an existing waypoint via mouse 2 or stylus on the touchscreen 3. Deleting a point at the end of the path is possible using a defined key on the keyboard 2 or a button on the touchscreen 3. To visualize to the RO whether the specified path is compatible with the vehicle’s kinematics, the center line is red-colored as soon as the curvature is outside the maximum steering angle of the vehicle. In order to understand the vehicle behavior generated by the RO input, the user can display the route calculated by the vehicle by using a defined key 2 or a button 3. After entering the path, the RO can set the vehicle in motion by accelerating or decelerating it using defined keys on the keyboard 2 or physical buttons on the touchscreen 3. In comparison, in Direct Control 1 the vehicle is controlled via steering angle inputs and acceleration or braking via buttons on the steering wheel.
@@ IV Results
Regarding the implementation of the interaction concept with separate path and velocity input in 2 and 3, it was noted that the input of the trajectory was intuitive, regardless of the input device. The adaptation of points and the simulated path calculated by the vehicle were also mentioned positively. However, one subject criticized the curve radius setting. Another criticism was that the line generated by the points was not intuitive, as the vehicle tended to cut the curves. Opinions were divided on planning waypoints and monitoring at the same time, as it would not be possible to concentrate on both at the same time. Moreover, test subjects criticized the orientation of the map, as it rotated in line with the vehicle’s movements. Also, the area to be driven through could not be seen at a glance on the map. Nevertheless, the map was described as particularly useful when the video quality was poor. Accordingly, the participants wanted a fixed map that can be moved and zoomed by the RO. Considering the different input devices, it was mentioned that the keyboard would be more intuitive to set the velocity but the stylus and tablet were found to be more comfortable by another test person. In terms of screen layout, it was noted that in the TS variant, the focus is on the map rather than the camera image. To avoid switching views, the information should be displayed on one screen or as an overlay where possible. A warning system or collision information was requested as a further overlay to the camera image.

### v2

@@ II-A Taxonomy
A teleoperation system comprises three components [10], as shown in Fig. 1:
@@ II-C Separate Path and Velocity Specification
Initial research into control via waypoints sent to a real vehicle was conducted by Kay [32] in a study with 19 participants. Vehicle velocity was controlled by a safety driver in the vehicle. The aim of the study was to investigate the functionality and performance of the waypoint specification under the aspect of latency and limited image data by measuring Key Performance Indicators like retries, total time or mean number of sent points. Effort in improving the RO workload is missing.
@@ III-A Design
To decide on an HMI design for separate path and velocity input, functionalities were collected in a morphological box and discussed with experts from the Institute of Automotive Technology at the Technical University of Munich. The discussion resulted in an HMI design that contains the components shown in Fig. 3 and the attached video link [^1] Video of the HMI: https://youtu.be/_e8sBJqupbg.
To enable the RO to set a valid path, the vehicle’s surroundings are represented on two screens showing a camera image and a map where the waypoints are set. Thereby, the monitor and laptop only serve as display devices. First, the RO specifies individual waypoints on the map by clicking the mouse in HMI variant 2 or using the stylus on the touchscreen in 3. The waypoints are automatically combined in the clicked sequence to form a driving path. During the interaction, a new waypoint is already selected, so the RO only has to set the point with one click. Furthermore, the user can edit the path by clicking, holding and moving an existing waypoint via mouse 2 or stylus on the touchscreen 3. Deleting a point at the end of the path is possible using a defined key on the keyboard 2 or a button on the touchscreen 3. To visualize to the RO whether the specified path is compatible with the vehicle’s kinematics, the center line is red-colored as soon as the curvature is outside the maximum steering angle of the vehicle. In order to understand the vehicle behavior generated by the RO input, the user can display the route calculated by the vehicle in red by using a defined key 2 or a button 3. After entering the path, the RO can set the vehicle in motion by accelerating or decelerating it using defined keys on the keyboard 2 or physical buttons on the touchscreen 3. In comparison, in Direct Control 1 the vehicle is controlled via steering angle inputs and acceleration or braking via buttons on the steering wheel.
@@ IV Results
Regarding the implementation of the interaction concept with separate path and velocity input in 2 and 3, it was noted that the input of the trajectory was intuitive, regardless of the input device. The adaptation of points and the simulated path calculated by the vehicle were also mentioned positively. However, one subject criticized the curve radius setting. Another criticism was that the line generated by the points was not intuitive, as the vehicle tended to cut the curves. Opinions were divided on planning waypoints and monitoring at the same time, as it would not be possible to concentrate on both. Moreover, test subjects criticized the orientation of the map, as it rotated in line with the vehicle’s movements. Also, the area to be driven through could not be seen at a glance on the map. Nevertheless, the map was described as particularly useful when the video quality was poor. Accordingly, the participants wanted a fixed map that can be moved and zoomed by the RO. Considering the different input devices, it was mentioned that the keyboard would be more intuitive to set the velocity but the stylus and tablet were found to be more comfortable by another test person. In terms of screen layout, it was noted that in the TS variant, the focus is on the map rather than the camera image. To avoid switching views, the information should be displayed on one screen or as an overlay where possible. A warning system or collision information was requested as a further overlay to the camera image.